# Heavy-quark diffusion in 2+1D and Glasma-like plasmas: evidence of a transport peak

### v1

@@ I Introduction
In the initial stages of heavy ion collisions, an exotic state of QCD matter, the quark-gluon plasma (QGP), is formed Busza:2018rrf . Despite intense studies, understanding the properties in different stages of its evolution, particularly the pre-equilibrium matter, remains a significant challenge in theoretical physics. These early stages may exhibit unique and non-perturbative properties that are not easily accessible through direct observation or computational techniques. As a result, open questions persist regarding the behavior of the QGP during its approach toward a hydrodynamic evolution.
@@
This work is structured as follows: In Sec. II, we review the theoretical background, our setup, and relevant quantities computed in this paper. We introduce the heavy-quark diffusion coefficient $\kappa(t,\Delta t)$ in Sec. III and discuss its properties such as its self-similarity and (universal) time-$\Delta t$ evolution. In Sec. IV, we reconstruct this observable using correlation functions in the frequency domain. In Sec. V, we systematically manipulate the excitation spectrum and study the impact on the heavy-quark diffusion coefficients. We discuss the implications of our findings for heavy-ion collisions and especially for the Glasma in Sec. VI, before concluding in Sec. VII. Further details of the HTL formalism used in the spectral reconstruction procedure and our results on self-similar properties can be found in Appendices A and B.
@@ II.2 Spectral and statistical correlation functions
Finally, we comment on the gauge (in-)variance of these correlation functions. Due to the employed temporal gauge $A_{0}=0$, the correlator $\langle EE\rangle(t,t^{\prime})$, and as a consequence also $\langle EE\rangle(t,\omega)$, etc., can be written in a manifestly gauge-invariant way Ipp:2020mjc ,
@@ III.3 Numerical computation of $\boldsymbol{\kappa(t,\Delta t)}$
Fig. 1: Heavy-quark diffusion coefficient $2\kappa$ for the 2D system (top), the gluonic sector of the Glasma-like system (center) and its scalar contribution $\kappa_{z}$ (bottom) as functions of $\Delta t$ at $Qt=500$. The insets depict early times that have been partially cut in the main panels. We also compare different expressions $\kappa^{\lessgtr}$, $\kappa_{\mathrm{def}}$ and $\kappa_{\mathrm{rec}}$ given by Eqs. (16)-(18) and the reconstructed curves as discussed in Sec. IV.
@@ IV.1 Properties of the spectral and statistical functions
$\frac{\dot{\rho}_{\alpha}(t,\omega,p)}{\dot{\rho}_{\alpha}(t,t,p)}\approx h(\omega,p)+\frac{\omega\ \rho_{\alpha}^{\mathrm{Landau}}(t,\omega,p)}{\dot{\rho}_{\alpha}(t,t,p))}$ (34)
@@ V.2 Implications from broad and narrow excitations
A more substantial effect appears in the scalar sector $\kappa_{\mathrm{z}}$, where the excitation peaks at low momenta are narrow similar to the narrow peaks in 3+1D Boguslavski:2018beu . An increase in the peak width at low momenta generates a noticeable offset at $\omega=0$ and leads to an evolution of $\kappa_{z}$ that closely resembles $2\kappa$ of the gluonic sector. This includes a shift of the late $\Delta t$ values upwards with strongly damped late-time oscillations. Reducing the damping rate of the already narrow scalar excitation peaks effectively improves the agreement with the original $\kappa_{z}$. This indicates that we may have overestimated the width of the scalar excitations at low momenta, which could be a consequence of the Fourier transform in Boguslavski:2021buh as mentioned in footnote 5. We conclude that the distinct oscillations around zero, and thus the non-diffusive character of $\kappa_{\mathrm{z}}$, seem to originate from the narrow quasi-particle-like excitations of the scalar sector.
@@ V.3 Evidence of an infrared enhancement of the scalar sector
Throughout this work, the scalar correlator $\kappa_{\mathrm{z}}(t,\Delta t)$ in the Glasma-like theory has been found to exhibit damped oscillations approximately around zero as a function of $\Delta t$, which is apparent in the lower panel of Fig. 1. We argued in the previous section V.2 that these oscillations are connected to the narrow scalar excitation peaks of $\langle EE\rangle_{z}(t,\omega,p\lesssim m_{D})$. We show here that there is another exciting factor contributing to the evolution of the scalar heavy-quark diffusion coefficient: an excess of soft gluons in the equal-time correlation function $\langle EE\rangle_{z}(t,t,p)$ at low momenta as compared to $\langle EE\rangle_{T}(t,t,p)$, which was observed in Ref. Boguslavski:2019fsb . This infrared (IR) enhancement is shown in the upper panel of Fig. 10 where we show both equal-time correlators as functions of momentum. While they coincide and rapidly fall off towards 0 for higher momenta, the scalar correlator is strongly enhanced at small momenta. We aim to demonstrate that it affects $\kappa_{z}(t,\Delta t)$ by repeating its spectral reconstruction, but this time with a suppressed equal-time correlator at low momenta.
@@ VI Consequences for heavy-ion collisions
The heavy-quark diffusion coefficient has been studied extensively in Refs. Carrington:2020sww ; Boguslavski:2020tqz ; Khowal:2021zoo ; Avramescu:2023qvv ; Das:2015aga ; Mrowczynski:2017kso ; Ipp:2020mjc ; Pandey:2023dzz in the Glasma and similar models. We will focus on a comparison with the recent study Avramescu:2023qvv . Very similar to that work, we find a pronounced peak at early times $Q\Delta t\lesssim 1$ of the evolution of $\kappa$ for the gluonic (their transverse) as well as of $\kappa_{z}$ for the scalar (their longitudinal) contributions, as we showed in the insets of Fig. 1. They indeed grow linearly in $\Delta t$ and the slope is determined by the energy density of the respective sector, as explained in Eq. (23) and in Ref. Boguslavski:2020tqz . This implies that the heavy-quark momentum broadens as $\langle p^{2}\rangle\propto\Delta t^{2}$ initially. In our case, we have roughly the same energy density in the gluonic and scalar contributions, and thus, this early-time evolution is quantitatively similar for $2\kappa$ and $\kappa_{z}$ (see insets of Fig. 1). In contrast, in the Glasma one finds $\kappa_{z}>2\kappa$ initially, which may hint at a slightly larger energy density in the scalar sector than for the gluonic components.
@@ VII Conclusion
  • soft gluons being short-lived while soft scalar excitations having extended lifetimes, which are visible as broad and narrow excitation peaks in gluonic $\langle EE\rangle$ and scalar $\langle EE\rangle_{z}$ for $p\lesssim m_{D}$,
@@
Another exciting prospect would be to compute $\kappa(t,\Delta t)$ in full quantum field theory. While conventional lattice gauge theory Brambilla:2020siz ; Banerjee:2022gen ; Brambilla:2022xbd ; Altenkort:2023oms ; Altenkort:2023eav ; Laine:2009dd can only extract its late-time value $\kappa_{\infty}=\lim_{\Delta t\to\infty}\kappa(t,\Delta t)$ in thermal equilibrium, the use of complex Langevin for real-time correlation functions Boguslavski:2022dee ; Boguslavski:2023unu may enable a new avenue towards extracting the entire time evolution from first principles.
@@ Appendix A HTL expressions
$+2\pi Z_{\alpha}(p)\left[\delta(\omega-\omega_{\alpha}^{\mathrm{HTL}}(p))-(w\mapsto-w)\right].$
@@ Appendix B Scaling behavior of fit parameters
These parameters are well described by a scaling law $\sim Q(Qt)^{-\sigma}$. We have summarized the resulting scaling exponents from a least square fitting procedure through our numerical data in Table 1 for the considered theories and both gluonic and scalar sectors. The extracted parameters are shown with the corresponding fitted power laws as functions of time in Fig. 12. While they generally agree well with the power law fits, slight deviations between the data and the curves are still visible, especially for the 2D theory. This is not surprising since the 2D correlators exhibit weaker initial oscillations and larger error bars than their gluonic 2D+sc counterparts as visible in the upper panel of Fig. 2. Approximating them as damped harmonic oscillators is, therefore, more difficult and prone to numerical inaccuracies, which is reflected in the 2D scaling curve fits not being as accurate as for the Glasma-like system. The top row in Fig. 12 shows the amplitude $A_{\mathrm{fit}}$ of $\kappa_{\mathrm{fit}}(t,\Delta t)$ while the center and bottom rows show the frequency $\omega_{\mathrm{fit}}$ and decay rate $\gamma_{\mathrm{fit}}$.

### v2

@@ I Introduction
In the initial stages of heavy-ion collisions, an exotic state of QCD matter, the quark-gluon plasma (QGP), is formed Busza:2018rrf . Despite intense studies, understanding the properties in different stages of its evolution, particularly the pre-equilibrium matter, remains a significant challenge in theoretical physics. These early stages may exhibit unique and non-perturbative properties that are not easily accessible through direct observation or computational techniques. As a result, open questions persist regarding the behavior of the QGP during its approach toward a hydrodynamic evolution.
@@
This work is structured as follows: In Sec. II, we review the theoretical background, our setup, and relevant quantities computed in this paper. We introduce the heavy-quark diffusion coefficient $\kappa(t,\Delta t)$ in Sec. III and discuss its properties such as its self-similarity and (universal) time-$\Delta t$ evolution. In Sec. IV, we reconstruct this observable using correlation functions in the frequency domain. In Sec. V, we systematically manipulate the excitation spectrum and study the impact on the heavy-quark diffusion coefficients. We discuss the implications of our findings for heavy-ion collisions and especially for the Glasma in Sec. VI, before concluding in Sec. VII. Further details of the HTL formalism used in the spectral reconstruction procedure and our results on self-similar properties can be found in Appendices A and B. The data for the computations performed in this work, including scripts for the reproduction of the figures, are available in backfried_2024_14039178 .
@@ II.2 Spectral and statistical correlation functions
Finally, we comment on the gauge (in-)variance of these correlation functions. Due to the employed temporal gauge $A_{0}=0$, the correlator $\langle EE\rangle(t,t^{\prime})$, and as a consequence also $\langle EE\rangle(t,\omega)$, can be written in a manifestly gauge-invariant way Ipp:2020mjc ,
@@ III.1 Heavy-quark diffusion
More precisely, the expressions in Eqs. (16) - (18) correspond rather to the quarkonium transport coefficient, which is in general different from the heavy quark momentum diffusion coefficient Scheihing-Hitschfeld:2022xqx . However, it has recently been shown in classical-statistical simulations of the Glasma Avramescu:2023qvv and of 3+1 dimensional gluonic plasmas Pandey:2023dzz that these definitions coincide with the momentum diffusion of heavy quarks.
@@ III.3 Numerical computation of $\boldsymbol{\kappa(t,\Delta t)}$
Fig. 1: Heavy-quark diffusion coefficient $2\kappa$ for the 2D system (top), the gluonic sector of the Glasma-like system (center) and its scalar contribution $\kappa_{z}$ (bottom) as functions of $\Delta t$ at $Qt=500$. The insets depict early times that have been partially cut in the main panels. We also compare different expressions $\kappa^{\lessgtr}$ and $\kappa_{\mathrm{def}}$ given by Eqs. (16)-(18) and the reconstructed curves $\kappa_{\mathrm{rec}}$ that will be discussed in Sec. IV.
@@ IV.1 Properties of the spectral and statistical functions
$\frac{\dot{\rho}_{\alpha}(t,\omega,p)}{\dot{\rho}_{\alpha}(t,t,p)}\approx h(\omega,p)+\frac{\omega\ \rho_{\alpha}^{\mathrm{Landau}}(t,\omega,p)}{\dot{\rho}_{\alpha}(t,t,p)}$ (34)
@@ V.2 Implications from broad and narrow excitations
A more substantial effect appears in the scalar sector $\kappa_{z}$, where the excitation peaks at low momenta are narrow similar to the narrow peaks in 3+1D Boguslavski:2018beu . An increase in the peak width at low momenta generates a noticeable offset at $\omega=0$ and leads to an evolution of $\kappa_{z}$ that closely resembles $2\kappa$ of the gluonic sector. This includes a shift of the late $\Delta t$ values upward with strongly damped late-time oscillations. Reducing the damping rate of the already narrow scalar excitation peaks effectively improves the agreement with the original $\kappa_{z}$. This indicates that we may have overestimated the width of the scalar excitations at low momenta, which could be a consequence of the Fourier transform in Boguslavski:2021buh as mentioned in footnote 5. We conclude that the distinct oscillations around zero, and thus the non-diffusive character of $\kappa_{z}$, seem to originate from the narrow quasi-particle-like excitations of the scalar sector.
@@ V.3 Evidence of an infrared enhancement of the scalar sector
Throughout this work, the scalar correlator $\kappa_{z}(t,\Delta t)$ in the Glasma-like theory has been found to exhibit damped oscillations approximately around zero as a function of $\Delta t$, which is apparent in the lower panel of Fig. 1. We argued in the previous section V.2 that these oscillations are connected to the narrow scalar excitation peaks of $\langle EE\rangle_{z}(t,\omega,p\lesssim m_{D})$. We show here that there is another exciting factor contributing to the evolution of the scalar heavy-quark diffusion coefficient: an excess of soft gluons in the equal-time correlation function $\langle EE\rangle_{z}(t,t,p)$ at low momenta as compared to $\langle EE\rangle_{T}(t,t,p)$, which was observed in Ref. Boguslavski:2019fsb . This infrared (IR) enhancement is shown in the upper panel of Fig. 10 where we show both equal-time correlators as functions of momentum. While they coincide and rapidly fall off toward 0 for higher momenta, the scalar correlator is strongly enhanced at small momenta. We aim to demonstrate that it affects $\kappa_{z}(t,\Delta t)$ by repeating its spectral reconstruction, but this time with a suppressed equal-time correlator at low momenta.
@@ VI Consequences for heavy-ion collisions
The heavy-quark diffusion coefficient has been studied extensively in Refs. Carrington:2020sww ; Boguslavski:2020tqz ; Khowal:2021zoo ; Avramescu:2023qvv ; Das:2015aga ; Mrowczynski:2017kso ; Ipp:2020mjc ; Pandey:2023dzz in the Glasma and similar models. We will focus on a comparison with the recent study Avramescu:2023qvv . Very similar to that work, we find a pronounced peak at early times $Q\Delta t\lesssim 1$ of the evolution of $\kappa$ for the gluonic (their transverse) as well as of $\kappa_{z}$ for the scalar (their longitudinal) contributions, as we show in the insets of Fig. 1. They indeed grow linearly in $\Delta t$ and the slope is determined by the energy density of the respective sector, as explained in Eq. (23) and in Ref. Boguslavski:2020tqz . This implies that the heavy-quark momentum broadens as $\langle p^{2}\rangle\propto\Delta t^{2}$ initially. In our case, we have roughly the same energy density in the gluonic and scalar contributions, and thus, this early-time evolution is quantitatively similar for $2\kappa$ and $\kappa_{z}$ (see insets of Fig. 1). In contrast, in the Glasma one finds $\kappa_{z}>2\kappa$ initially, which may hint at a slightly larger energy density in the scalar sector than for the gluonic components.
@@ VII Conclusion
  • soft gluons being short-lived while soft scalar excitations have extended lifetimes, which are visible as broad and narrow excitation peaks in gluonic $\langle EE\rangle$ and scalar $\langle EE\rangle_{z}$ for $p\lesssim m_{D}$,
@@
Another exciting prospect would be to compute $\kappa(t,\Delta t)$ in the full quantum field theory. While conventional lattice gauge theory Brambilla:2020siz ; Banerjee:2022gen ; Brambilla:2022xbd ; Altenkort:2023oms ; Altenkort:2023eav ; Laine:2009dd can only extract its late-time value $\kappa_{\infty}=\lim_{\Delta t\to\infty}\kappa(t,\Delta t)$ in thermal equilibrium, the use of complex Langevin for real-time correlation functions Boguslavski:2022dee ; Boguslavski:2023unu may enable a new avenue toward extracting the entire time evolution from first principles.
@@ Appendix A HTL expressions
$+2\pi Z_{\alpha}(p)\left[\delta(\omega-\omega_{\alpha}^{\mathrm{HTL}}(p))-(\omega\mapsto-\omega)\right].$
@@ Appendix B Scaling behavior of fit parameters
These parameters are well described by a scaling law $\sim Q(Qt)^{-\sigma}$. We have summarized the resulting scaling exponents from a least square fitting procedure through our numerical data in Table 1 for the considered theories and both gluonic and scalar sectors. The extracted parameters are shown with the corresponding fitted power laws as functions of time in Fig. 12. While they generally agree well with the power law fits, slight deviations between the data and the curves are still visible, especially for the 2D theory. This is not surprising since the 2D correlators exhibit weaker initial oscillations and larger error bars than their gluonic 2D+sc counterparts as is visible in the upper panel of Fig. 2. Approximating them as damped harmonic oscillators is, therefore, more difficult and prone to numerical inaccuracies, which is reflected in the 2D scaling curve fits not being as accurate as for the Glasma-like system. The top row in Fig. 12 shows the amplitude $A_{\mathrm{fit}}$ of $\kappa_{\mathrm{fit}}(t,\Delta t)$ while the center and bottom rows show the frequency $\omega_{\mathrm{fit}}$ and decay rate $\gamma_{\mathrm{fit}}$.